# A Fictional Q&A Dataset for Studying Memorization and Knowledge Acquisition

### v1

@@ 5 Experiments
Figure 8: Multiple choice accuracy with 4 choices as a function of optimization step for the Llama-3.1-8B model across fictional splits separating performance on the questions that were generated based on documents in the training set versus in the validation set for that split.
We observe that the the improvement in MCQ accuracy appears to be “leaky”. Figure 8 shows that the performance on the questions corresponding to held-out fictional scenarios and events is also elevated “(Val)”, albeit slightly less than for questions corresponding to scenarios that were trained on “(Train)” even though some facts in the Event Split are expected to have been wholly omitted from the training document pool. This indicates that it is not possible to cleanly differentiate whether the improvements we observe in MCQ Acc (or question conditional answer loss in Figure 6) are attribute-able to factual or stylistic memorization alone. However, it is clear that these improvements are caused by training on the fictional data and not other effects based on the lack of improvement for the “Base Webtext”-only control in Figure 7. All we can conclude is that the model’s improved ability to rank answer choices after training on the fictional documents in each of the experiments is based on some combination of distributional features and atomic knowledge it acquires during the finetuning process.
@@ 6 Discussion and Conclusion
We believe that our final results suggesting that the knowledge acquisition mechanism is “non-identifiabile”, and that the consistent result that training on the Fictsheets split offers the least amount of validation performance improvement in terms of loss or Q&A accuracy (Figures 5, 6, 7 and 8), are actually more compatible than they might seem at first glance. To humans, the clean, markdown-like structure of the fictsheets might appear to be the surface form of the factual information that is the easiest to extract generalizable knowledge from. However, LLMs acquire knowledge through different mechanisms than humans, relying on distributional features in the text more than anticipated or desired. We hypothesize that this set of results has implications for the conditions under which factual memorization is likely to occur, or possible at all.
While in Appendix A we discuss limitations and future applications in greater detail, we conclude with a few key remarks here. Constructing fully synthetic “cleanroom” data using LLMs is difficult. We design our prompts carefully to ensure the quality and diversity of the various components in our dataset but still observe a significant about of duplicate questions. The results in Figure 8 also suggest that the fictional documents might overlap to a larger degree than desired across seed events and across styles. While we elucidate interesting memorization vs. generalization behaviors through our experiments, more than anything, we hope that our results inspire the use of our dataset for studying topics we do not explore such as machine unlearning and privacy preserving training methods.
Acknowledgments and Disclosure of Funding
We gratefully acknowledge the contributions of Max Cembalest, who implemented an early version of our dataset generation pipeline; Christopher A. Choquette-Choo, Matthew Jagielski, and Avi Schwarzschild, who brainstormed and helped frame our initial research questions; Katherine Lee, who has provided invaluable feedback and insights throughout the project; and Sean McLeish, for his proofreading assistance.
Funding
@@
Appendix A Limitations and Future Applications
@@
We reformat the question and answer pairs from our pipeline into MCQs in a post-hoc manner, and this introduces artifacts (see Section C.6 for details on construction). We suspect that trivially easy to eliminate alternate answers cause the model to achieve base accuracies better than chance without actually training on the questions. We also suspect that, as a side effect of our ranking criteria during construction, multiple answer paraphrases or plausible alternative answers to vague questions end up in the alternates list, potentially upper-bounding the best case accuracy. This could generally be ameliorated in future work in various ways including creating the alternate lists for these MCQs from scratch during initial question generation, using more complex methods for creating alternate answer lists, and even via human curation or annotation of MCQs for feasibility and difficulty.
@@
Appendix B Extended Related Work
@@
B.1 The Impact of Repetition on Memorization and Model Capability
Large language models have been shown to memorize parts of their pretraining data in many different settings. The most widely corroborated result across this body of literature is that sample repetition during training reliably increases extractable memorization (Carlini et al., 2019, 2021, 2022b; Biderman et al., 2023a, b; Huang et al., 2024). Training data repetition, and the factual memorization it often entails, also impacts model performance in complex ways. Entity repetition has been shown to correlate with knowledge intensive benchmark performance (Kandpal et al., 2023), however, too much repetition also can adversely effect model performance (Muennighoff et al., 2023) and carefully deduplicating a pretraining corpus has been shown to simultaneously reduce memorization rates and often improve overall model performance (Kandpal et al., 2022; Lee et al., 2022; Tirumala et al., 2023). The literature also consistently shows that larger models exhibit larger rates of memorization and can exhibit this behavior after fewer repetitions of the training data (Carlini et al., 2022b; Duan et al., 2024; Singh et al., 2024).
B.2 Training Data Extraction and MIAs
@@ B.4 Forgetting and Unlearning
B.5 Generating Synthetic Datasets with LLMs
@@
B.6 Knowledge Acquisition
Since the the advent of GPT-3, users have become accustomed to the fact that LLMs absorb massive amounts of information about the world through their web scale pretraining process and are then able to demonstrate this knowledge in response to user prompts and task descriptions (Brown et al., 2020). The entire field of instruction finetuning, and a significant amount of all other post-training research, has been focused on increasing the ease with which users can unlock the knowledge intensive capabilities of base pretrained models even further. However, the literature on exactly how language models perceive, store, and do/do not demonstrate knowledge is much less mature. Seminal work by Kandpal et al. (2023) showed a clean relationship between entity co-occurrence in a training corpus and test time associative ability between those entities. More recently, the “Physics of LLMs” line of work (Allen-Zhu and Li, 2023a, b, 2024) studies small language models, trained on limited, but highly controlled datasets to try and uncover causal mechanisms for knowledge storage and production in LLMs.  Berglund et al. (2023) specifically studied the asymmetry in how LLMs generalize across declarative and interrogative forms of the same knowledge using synthetic data (eg. “A is B” vs “B is A” or “B was?”) and in the past year Chang et al. (2024); Zucchet et al. (2025) have both studied knowledge acquisition from the perspective of dynamics and training hyperparameters using synthetic data.
Appendix C Additional Generation and Annotation Pipeline Details
C.1 Dataset Release Schema

### v2

@@ 4.4 Reconstruction of Fictional Facts via MCQ Testing
Figure 8: MCQ accuracy over 4 choices as a function of optimization step for the Llama-3.1-8B model with performance separated for questions corresponding to documents in the training set vs. the validation set for each style split.
We observe that the the improvement in MCQ accuracy appears to be “leaky”. Figure 8 shows that the performance on the questions corresponding to held-out fictional scenarios and events is also elevated “(Val)”, albeit slightly less than for questions corresponding to scenarios that were trained on “(Train)”. This performance elevation is observed even though some facts underlying the MCQ’s corresponding to the Event Split’s validation set are expected to have been wholly omitted from the training document pool. This indicates that it is not possible to cleanly differentiate whether the improvements we observe in MCQ Acc. (or question-conditional answer loss in Figure 6) are attribute-able to either purely factual or purely stylistic memorization.
What is clear is that these improvements are caused by training on the fictional data and not other effects as evidenced by the lack of improvement for the “Base Webtext”-only control (Figure 7). All we can conclude is that the models’ improved MCQ accuracy after training on the fictional documents in each of the experiments is based on some combination of distributional and atomic learning. To sanity check these findings, we analyze the Event Split MCQs in detail in Appendix B.
@@ 5 Discussion and Conclusion
We believe that the observations that the knowledge acquisition mechanism under standard finetuning is not purely “atomic” and that training on the Fictsheets split offers the least amount of validation performance improvement (Figures 5, 6, 7 and 8), are actually more compatible findings than they might seem at first glance. To a human, the clean, markdown-like structure of the fictsheets might appear to be the surface form of the factual information that is the easiest to extract generalizable knowledge from. However, LLMs acquire knowledge through different mechanisms than humans, perhaps relying on distributional features in the text more than anticipated or desired. We hypothesize that this set of results has implications for the conditions under which factual memorization under limited repetition is likely to occur, or possible at all.
While in Appendix A we discuss limitations and future applications in greater detail, we conclude with a few key remarks here. Constructing fully synthetic “cleanroom” data using LLMs is difficult. We design our prompts carefully to ensure the quality and diversity of the various components in our dataset but still observe a significant number of duplicate questions. The results in Figure 8 also suggest that the fictional documents might overlap to a larger degree than desired across seed events and across styles. While we elucidate interesting memorization vs. generalization behaviors through our experiments, more than anything, we hope that our results inspire the use of our dataset for studying topics we do not explore such as machine unlearning and privacy preserving training methods.
Acknowledgements
We gratefully acknowledge the contributions of Max Cembalest, who implemented an early version of our dataset generation pipeline; Christopher A. Choquette-Choo, Matthew Jagielski, and Avi Schwarzschild, who brainstormed and helped frame our initial research questions; Katherine Lee, who provided invaluable feedback and insights throughout the project; and Sean McLeish, for his proofreading.
@@ Acknowledgements
Appendix A Extended Discussion of Limitations and Future Applications
@@
We reformat the question and answer pairs from our pipeline into MCQs in a post-hoc manner, and this introduces artifacts (see Section D.6 for details on construction). We observe that trivially easy to eliminate alternate answers probably cause the model to achieve base accuracies better than chance without actually training on the questions. We also see that, as a side effect of our ranking criteria during construction, multiple answer paraphrases or plausible alternative answers to vague questions end up in the alternates list, potentially upper-bounding the best case accuracy. We discuss this issue in the context of the Event Split MCQs in Appendix B and explore an alternate method for MCQ generation in Section D.6.1.
These issues could be ameliorated in future work in various ways such as by creating the alternate lists for these MCQs from scratch during initial question generation (a la Section D.6.1) and annotating and curating MCQs by hand for feasibility and difficulty. We note that constructing multiple choice tests that avoid these issues is a rich area of study. Prior work has found that models can often be “right for the wrong reasons” which is one way of summarizing potential confounders in our MCQ setup (McCoy et al., 2019).
@@
Appendix B Extended Analysis of MCQs and Figure 8 Results
One goal underlying our dataset construction process (eg. the “Event Split”) is to make it possible to identify whether or not models finetuned on our data are more likely learning factual relationships atomically or via general distributional properties of the fictional data. In theory, this distinction can be made by checking whether they improve on only the questions associated with training documents, or whether they also improve on questions associated with the validation documents. Even from just Figure 4 (right), it is already clear that some sort of distributional transfer learning is happening across “Event Split” lines, so seeing some amount of similar transfer between the associated training and validation MCQ examples in Figure 8 is also expected. That being said, whether or not a specific model or training algorithm appears to learn more atomically or more distributionally is itself the actual empirical question at hand in these experiments. The goal in running these experiments using our dataset is to simply report the observations of what appears to be happening in a standard training setting, using our data as the probe, rather than to design some training recipe that targets a specific outcome.
With that as context, in this section we perform some additional analyses of the data behind Figure 8. In particular, we analyze the sample-wise performance of the Llama-3.1-8B model on the Event Split’s training MCQs and the validation MCQs and compare successes and failures at step 0 of training to those at the final step of training. The goal of this analysis is to try and discern whether there are any particular defining characteristics of the validation question subset that the Llama-3.1-8B model “learned” over the course of training, and whether there is any specific distributional link between the training questions the model learns during training, versus the validation questions that it simultaneously improved on.
B.0.1 A note on duplicated answers from distinct questions
We note that our question filtering process deduplicated the questions only on the question text, and not the answer itself. As an example, from two different documents based on the same seed event, we can see that two questions were produced that both have the answer “Lake Ypsilon”.
question_id: event_000_style_corporate_num_000_question_003
Question: Where was the first pilot test of the Ring of Silence Protocol conducted?
Answer: Lake Ypsilon
question_id: event_000_style_encyclopedia_num_000_question_002
Question: Where was the sound-absorbing moat first implemented?
Answer: Lake Ypsilon
However this is an artifact of various documents and their questions ending up referring to shared concepts in the same seed event even though the way in which the question is posed (i.e. the specific fact pattern) is slightly different. In the subsequent analyses we continue to present the questions as subselected for experiments in the main paper for continuity. However, during the analyses we also checked whether this additional deduplication by answer seemed to change any of the statistics meaningfully, but did not observe any evidence of that.
B.0.2 Visual inspection of the training questions versus validation questions
We first can take a look at some samples from the sets of training split questions the model improved on as well as samples from the set of validation questions the model improved on. One thing we observe is that there are indeed cases where there is more than one similar answer that could plausibly be. These cases can only be decided by specific formatting quirks aligned with the source fiction document such as using a full proper noun versus a shortened one. This at least suggests an explanation for the fact that accuracy plateaus around 50% even for the largest models at the end of training.
=== Sample 1 (Index 2832) ===
event_id: event_093
fiction_id: event_093_style_social_num_000
question_id: event_093_style_social_num_000_question_002
input: Question: Who supposedly left instructions for the steeple’s construction?
Answer:
target_hf: William Linton
target_idx: 2
topk_choices:
  Choice 0: Western and Eastern
  Choice 1: Raising the Brow: The Brow Society’s Legacy
  Choice 2: William Linton
  Choice 3: The Source
choice_ranking_init: [3, 2, 0, 1]
choice_ranking_final: [2, 3, 0, 1]
acc_init: 0.0
acc_final: 1.0
=== Sample 2 (Index 897) ===
event_id: event_029
fiction_id: event_029_style_news_num_003
question_id: event_029_style_news_num_003_question_001
input: Question: Who is the entrepreneur responsible for the AI-driven ice cream carts?
Answer:
target_hf: Lila Sorvino
target_idx: 3
topk_choices:
  Choice 0: Dr. Lila Harrington
  Choice 1: Lillian Abbott
  Choice 2: FlavorSync AI
  Choice 3: Lila Sorvino
choice_ranking_init: [1, 2, 0, 3]
choice_ranking_final: [3, 2, 0, 1]
acc_init: 0.0
acc_final: 1.0
  MCQ Examples: Random questions in the Event Split Train set for which the Llama-3.1-8B model initially answers incorrectly, but then answers correctly at the final step i.e. “learned”. Here we see that in Sample 1 (Index 2832) context clues about the entity type most likely eliminate all three other choices, but in Sample 2 (Index 897) there are three plausible answers all of the correct entity type.
The other consistent feature is that there are normally one or two distractor choices that most human readers could easily eliminate using question and context clues without needing to look at the fictional document itself. Similar to the aforementioned case of choices that are overly semantically similar, the presence of these obviously incorrect distractors suggests a partial explanation for why the models achieve better than chance ($1/\text{num choices}$) accuracy at step 0 in the main experiments. However, curiously, the Llama-3.1-8B model still often still ranks one of those “easily eliminable” choices in the top 2 indicating that these models’ likelihood assignments under the prompting and loss ranking process used to score the model on the MCQs do not always effectively eliminate obvious wrong answers in the way a good human test taker might. [^8] To address these issues, we explore an alternate methodology for generating MCQs that do not suffer from these issues as much. This version of the MCQ generation procedure and resulting assets are discussed in Section D.6.1.
=== Sample 1 (Index 2535) ===
event_id: event_084
fiction_id: event_084_style_corporate_num_000
question_id: event_084_style_corporate_num_000_question_004
input: Question: What did the diary become a symbol of?
Answer:
target_hf: shared suffering
target_idx: 1
topk_choices:
  Choice 0: shared suffering and humanity
  Choice 1: shared suffering
  Choice 2: vulnerabilities
  Choice 3: vivid sketches and poignant accounts
choice_ranking_init: [2, 1, 0, 3]
choice_ranking_final: [1, 2, 0, 3]
acc_init: 0.0
acc_final: 1.0
=== Sample 2 (Index $11$) ===
event_id: event_036
fiction_id: event_036_style_news_num_002
question_id: event_036_style_news_num_002_question_001
input: Question: Who discovered the Waterfall Whisper phenomenon?
Answer:
target_hf: Thomas Bright
target_idx: 3
topk_choices:
  Choice 0: tie
  Choice 1: Western and Eastern
  Choice 2: acoustic engineering and psychological principles
  Choice 3: Thomas Bright
choice_ranking_init: [0, 3, 1, 2]
choice_ranking_final: [3, 0, 1, 2]
acc_init: 0.0
acc_final: 1.0
  MCQ Examples: Random questions in the Event Split Validation set for which the Llama-3.1-8B model initially answers incorrectly, but then answers correctly at the final step, i.e. “learned”. In Sample 1 (Index 2535), three of the choices are quite semantically similar and thus equally plausible, whereas in Sample 2 (Index $11$) there is one answer that is more plausible than the others though a human might need to check against the source document to be absolutely sure.
Overall, we observe no telltale indicators of some characteristic difference between the questions the model gets correct after training in the training split or the validation split question sets but we acknowledge that our manual inspection is not exhaustive.
B.0.3 Error case analysis
For the Event Split’s two sets of training questions and validation questions we are able to analyze two different evaluation runs of the same model: one that was performed at the beginning of training, and another performed at the end of training. As shown in the examples above, we annotate the two result datasets with a boolean indicating whether the model answered each question correctly “initially” at step 0 and/or correctly at the “final” step of training. Then, we subset the data based on these two conditions (and a few others) to compute statistics that try and shed light on whether there are any interesting differences on how the model performance changed during the course of training.
Initial Acc Final Acc
Train Set 31.91% (633/1984) 45.31% (899/1984)
Validation Set 30.51% (321/1052) 43.44% (457/1052)
Table 1: The exact numbers underlying the start and end points of the (Train) and (Val) lines for the Event Split for the Llama-3.1-8B model in Figure 8. These are the same sets of questions and their correctness judgements that we use in the case-wise analysis in this section.
In Table 1 we present the aggregate statistics for the Llama-3.1-8B model on this set of MCQs separated by train and validation. Then, rather than show an extremely extensive table of all possible subsettings by taking the cross product of all conditions we could think of, we focus on just a few combinations of potential interest and present the results in Table 2. In particular, we use the terms “acquired”/“learned” to indicate subsetting for questions for which the model was initially incorrect, but then ended up answering correctly after the final step of training. We also note a particularly common spurious answer that ended up in a large number of the choice lists “Western and Eastern” (abbr. “W and E”), and also often ended up ranked as the model’s top choice at the first step of training. This is an artifact of our choice ranking process (see Section D.6) and could be improved in future work on similar pipelines.
The summary takeaway here is that none of the particular statistics we calculate are notably different between the training questions and the validation questions; all we see is that the training split yields slightly elevated correctness levels in all cases. So, under this lens of analysis, it appears that most of the information being learned by the model that helps it correctly answer new training questions it did not initially answer correctly, must also be similarly useful for helping it answer some additional validation questions more accurately. Or rather, as stated in Section 5, the models we train appear to be primarily learning distributionally, not atomically. One could interpret the slight additional improvement on the training split as the fraction of improvement attributable to more atomic knowledge acquisition, but our experiments do not prove this conclusively.
Case Train Split Val Split
Fraction of initially correct answers that still remained correct after training: 85.62% (542/633) 85.36% (274/321)
Fraction of final correct answers that were learned i.e. not initially correct: 39.71% (357/899) 40.04% (183/457)
Fraction of learned questions that also had grade_informed=1: 93.28% (333/357) 89.62% (164/183)
Fraction of learnable questions (initially incorrect) that were learned: 26.42% (357/1351) 25.03% (183/731)
Fraction of learned questions where natural_answer_in_fiction==1: 82.63% (295/357) 83.06% (152/183)
Fraction of questions that model could have learned but did not where natural_answer_in_fiction==1: 56.04% (557/994) 58.94% (323/548)
Fraction of learnable questions where natural_answer_in_fiction==1 that were actually learned: 34.62% (295/852) 32.00% (152/475)
Fraction of learnable questions where natural_answer_in_fiction==0 that were actually learned: 12.42% (62/499) 12.11% (31/256)
Fraction of samples with ‘W and E’ as a choice: 31.45% (624/1984) 32.41% (341/1052)
Fraction of learned questions that had ‘W and E’ as initial model choice but then demoted it: 26.05% (93/357) 29.51% (54/183)
Fraction of learnable questions that were learned ignoring W and E demotion cases: 20.99% (264/1258) 19.05% (129/677)
Table 2: Case wise error analysis of Event Split’s Train and Test sets of questions for the Llama-3.1-8B model.
B.0.4 Train and Validation Sample Distributional Overlap
Finally, to try and identify any additional links between the training questions and the validation questions that might explain the transfer learning indicated by the improvements in performance on validation questions, we extract all the named entities and whitespace separated tokens (words) from the fictional documents and the answers to the questions. Then, we compute the aggregate overlap between all entities or words from different combinations of the training fictions or answers and the validation fictions or answers. These results are presented in Tables 3, 4 and 5
The main hypothesis behind this analysis would be that the cases where we take the set of training questions that only became correct after training (directly “learned”) and the set of questions in the validation set where the model also happened to become correct after training (indirectly “transfer learned”) could show some elevated level of distributional overlap as compared to all possible subset comparisons. An observation like this would help explain why the model improved on those specific validation questions over others.
However, we find that the entity or word based Jaccard similarity (set intersection over set union) for that specific case (bolded row 5), no matter which combination of the fiction texts or answer texts with intersected, is never significantly higher than the other cases shown in other rows. While this could simply be too simple of a distributional comparison, it does indicate that the specific improvements in validation question performance cannot be explained by relatively high levels of simple word overlap alone. Rather, there is some collection of general distributional patterns that the model is learning from the training documents that is also useful in ranking validation answers more accurately.
Train Init Acc Train Final Acc Val Init Acc Val Final Acc Entity Jaccard Similarity Word Jaccard Similarity
0 0 0 0 0 0.062 0.367
1 0 0 0 1 0.049 0.299
2 0 0 1 0 0.022 0.183
3 0 0 1 1 0.055 0.332
4 0 1 0 0 0.066 0.382
5 0 1 0 1 0.058 0.347
6 0 1 1 0 0.032 0.234
7 0 1 1 1 0.064 0.366
8 1 0 0 0 0.053 0.33
9 1 0 0 1 0.063 0.355
10 1 0 1 0 0.044 0.301
11 1 0 1 1 0.06 0.35
12 1 1 0 0 0.064 0.388
13 1 1 0 1 0.056 0.335
14 1 1 1 0 0.028 0.216
15 1 1 1 1 0.062 0.363
Table 3: Case-wise analysis for the Event Split evaluation of the Llama-3.1-8B model, where similarities are computed between the Fictional Documents in Train split vs. Fictional Documents in Val split.
Train Init Acc Train Final Acc Val Init Acc Val Final Acc Entity Jaccard Similarity Word Jaccard Similarity
0 0 0 0 0 0.018 0.105
1 0 0 0 1 0.014 0.058
2 0 0 1 0 0 0.021
3 0 0 1 1 0.016 0.067
4 0 1 0 0 0.004 0.062
5 0 1 0 1 0.018 0.071
6 0 1 1 0 0 0.023
7 0 1 1 1 0.01 0.046
8 1 0 0 0 0 0.04
9 1 0 0 1 0 0.04
10 1 0 1 0 0 0.033
11 1 0 1 1 0.009 0.042
12 1 1 0 0 0.011 0.078
13 1 1 0 1 0.01 0.059
14 1 1 1 0 0 0.03
15 1 1 1 1 0.054 0.082
Table 4: Case-wise analysis for the Event Split evaluation of the Llama-3.1-8B model, where similarities are computed between the Fictional Answers in Train split vs. Fictional Answers in Val split.
Train Init Acc Train Final Acc Val Init Acc Val Final Acc Entity Jaccard Similarity Word Jaccard Similarity
0 0 0 0 0 0.002 0.022
1 0 0 0 1 0.001 0.007
2 0 0 1 0 0 0.003
3 0 0 1 1 0.005 0.01
4 0 1 0 0 0.003 0.029
5 0 1 0 1 0.002 0.009
6 0 1 1 0 0 0.004
7 0 1 1 1 0.009 0.013
8 1 0 0 0 0.007 0.044
9 1 0 0 1 0.003 0.015
10 1 0 1 0 0 0.006
11 1 0 1 1 0.013 0.02
12 1 1 0 0 0.003 0.027
13 1 1 0 1 0.001 0.009
14 1 1 1 0 0 0.004
15 1 1 1 1 0.007 0.012
Table 5: Case-wise analysis for the Event Split evaluation of the Llama-3.1-8B model, where similarities are computed between the Fictional Documents in Train split vs. Fictional Answers in Val split.
Appendix C Extended Related Work
@@
C.1 The Impact of Repetition on Memorization and Model Capability
Large language models have been shown to memorize parts of their pretraining data in many different settings. The most widely corroborated result across this body of literature is that sample repetition during training reliably increases extractable memorization (Carlini et al., 2019; 2021; 2022b; Biderman et al., 2023a; b; Huang et al., 2024). Training data repetition, and the factual memorization it often entails, also impacts model performance in complex ways. Entity repetition has been shown to correlate with knowledge intensive benchmark performance (Kandpal et al., 2023), however, too much repetition also can adversely effect model performance (Muennighoff et al., 2023) and carefully deduplicating a pretraining corpus has been shown to simultaneously reduce memorization rates and often improve overall model performance (Kandpal et al., 2022; Lee et al., 2022; Tirumala et al., 2023). The literature also consistently shows that larger models exhibit larger rates of memorization and can exhibit this behavior after fewer repetitions of the training data (Carlini et al., 2022b; Duan et al., 2024; Singh et al., 2024).
C.2 Training Data Extraction and MIAs
@@ C.4 Forgetting and Unlearning
C.5 Generating Synthetic Datasets with LLMs
@@
C.6 Knowledge Acquisition
Since the the advent of GPT-3, users have become accustomed to the fact that LLMs absorb massive amounts of information about the world through their web scale pretraining process and are then able to demonstrate this knowledge in response to user prompts and task descriptions (Brown et al., 2020). The entire field of instruction finetuning, and a significant amount of all other post-training research, has been focused on increasing the ease with which users can unlock the knowledge intensive capabilities of base pretrained models even further. However, the literature on exactly how language models perceive, store, and do/do not demonstrate knowledge is much less mature. Seminal work by Kandpal et al. (2023) showed a clean relationship between entity co-occurrence in a training corpus and test time associative ability between those entities. More recently, the “Physics of LLMs” line of work (Allen-Zhu and Li, 2023a; b; 2024) studies small language models, trained on limited, but highly controlled datasets to try and uncover causal mechanisms for knowledge storage and production in LLMs.  Berglund et al. (2023) specifically studied the asymmetry in how LLMs generalize across declarative and interrogative forms of the same knowledge using synthetic data (eg. “A is B” vs “B is A” or “B was?”) and in the past year Chang et al. (2024); Zucchet et al. (2025) have both studied knowledge acquisition from the perspective of dynamics and training hyperparameters using synthetic data.
Appendix D Additional Generation and Annotation Pipeline Details
D.1 Dataset Release Schema